09. A box of mass 1 Kg slides with initial velocity 6 m/s on a frictionless surface and collides inelastically with an identical rest box. The boxes stick together and move after the collision. What is the change of kinetic energy? A)-91 B)18] C)3] D) -15] E) 5
Answer #1

m1= 1kg, u1= 6m/s  m2= 1 kg, u2= 0

Apply law of conservation of momentum

m1u1 + m2u2 = (m1+m2)v

1× 6 + 1× 0 = 2×v

v = 3m/s

Change in kinetic energy

Final energy - initial energy= 1/2 (m1+m2) v^2 - 1/2m1u1^2= 1/2× 2 ×3 ×3 - 1/2× 1× 6×6= 9- 18 = -9 J

So answer is option A